# SystolicAttention: Fusing FlashAttention within a Single Systolic Array

### v2

@@ 3.5. Overlapping Computations with SystolicAttention
During the first matrix multiplication, the head dimension $d$ of both $Q$ and $K$ is mapped to $N\_ROWS$, while the sequence length $B_{r}$ of $Q$ is mapped to $N\_COLS$. After computing $S$, we re-stream it from the top of the array, mapping the $B_{c}$ dimension of $S$ to $N\_ROWS$. A sequence of element-wise operations is then performed in-place cycle by cycle to transform $S$ into $N$ and then into $P$.
For the second matrix multiplication, $O=P\cdot V$, the $B_{c}$ dimension of $V$ is mapped to $N\_ROWS$ and contracted with the stationary matrix $P$ inside the systolic array.
@@
Figure 6 illustrates the execution of the entire inner loop. The fifth operation, $a=old\_m-new\_m$, passes down the difference between the previous and current maximum values from comparison units to the accumulator. This operation runs in parallel with the constant multiplication $\frac{\log_{2}e}{\sqrt{d}}\cdot N$, as the downward data path is idle during the in-place constant multiplication.
Overall, each iteration takes only $102\cdot N\_COLS+3\cdot N\_ROWS+$ 10 cycles, meaning that on an $N\times N$ systolic array, it takes just $105N+$ 10 cycles to process a $N\times N$ FlashAttention tile. By contrast, on a naive $N\times N$ systolic array, the two independent matrix multiplications alone may require up to $28N-$ 2 cycles due to preloading and synchronization overheads discussed in Section 2.1.
The re-scaling operation in Line 21 of Algorithm 1 is only performed once per outer loop. It is executed using the accumulator after the second matrix multiplication completes. In our implementation, this re-scaling step takes $202N+$ 20 cycles, which is negligible compared to the inner loop execution time.
@@ 4.1. Microarchitecture Overview
The instruction decoder classifies instructions into three types: load, store, and compute. Instructions of different classes are executed asynchronously, while instructions of the same class execute in order but may be interleaved.
@@ 5.2. Python API for FSA Instructions
Each FSA instruction is exposed through a corresponding Python API. These APIs are designed to be type-safe, ensuring correct usage of tensor types. The complete set of supported APIs is shown in Listing 1.
@@ 5.3. JIT Kernel Compiler
The FlashAttention kernel implemented in FSA using the JIT compiler is shown in Listing 2.
@@
In the example above, the host loads input tensors into device memory using DRAMSim2 (Rosenfeld et al., 2011), dispatches the compiled instructions to the FSA device, and retrieves the output tensors upon completion. The output is then converted to NumPy (Harris et al., 2020) arrays for post-processing.
Since FSA currently does not support matrix transpose operations in hardware, the input matrix $V$ must be transposed in advance, and the resulting output matrix $O$ is returned in transposed form.
@@ 6.1. FlashAttention Performance
To clarify, we use the term FLOPs to denote the number of floating-point operations, and FLOPs/s to indicate the number of floating-point operations performed per second. Although all accelerators share the same systolic array size, their absolute performance still varies due to differences in clock frequency, which in turn depend on the specific node technology employed. Therefore, we simulate FSA at 1 GHz, a target that is easily achievable in 16 nm technology, as discussed in Section 6.3, and compare matrix multiplication FLOPs/s utilization rather than wall-clock execution time.
@@
Systolic array size $128\times 128$ $128\times 128$ $128\times 128$
@@
Frequency 1.5GHz1 2.8GHz 1GHz
Memory Bandwidth 819GB/s 820GB/s 512GB/s
@@
The number of floating-point operations in a FlashAttention inner loop is given by:
[…]
where each matrix multiplication contributes $2\times B_{r}\times B_{c}\times d$. The total number of operations is:
$\text{FLOPs}=\frac{\text{SeqLen}}{B_{r}}\times\text{FLOPs/InnerLoop}\times\frac{\text{SeqLen}}{B_{c}}=4\times\text{SeqLen}^{2}\times d$
This matches the formula used in FlashAttention (Dao, 2023).
@@
[…]
@@
[…]
@@
  • jax.pallas.ops.tpu.flash_attention for TPUv5e (JAX Developers, [n. d.]).
@@
The experimental results are shown in Figure 10. Benefiting from intensively overlapped computations, FSA achieves average FLOPs/s utilization that is 1.77 times and 4.83 times higher compared to TPUv5e and AWS NeuronCore-v2, respectively. Moreover, the results confirm that our RTL implementation of FSA closely aligns with the theoretical performance outlined in Section 3.5, validating the correctness of our RTL design.
@@ 6.2.1. Error Analysis of Piecewise Linear Approximation
Figure 11 shows the mean absolute error (MAE) and mean relative error (MRE) of the piecewise linear approximation of the exp2 function, plotted against the number of segments used in the approximation.
The MAE decreases significantly as the number of segments increases, while the MRE remains relatively stable. This occurs because the breakpoints in the approximation are uniformly spaced, whereas the fp16 representation is logarithmic in nature, making relative error more sensitive for values close to zero.
We choose to use 8 segments in our FSA implementation, which achieves a MAE of 0.00014 and an MRE of 0.02728. Although non-uniformly spaced breakpoints could further reduce the MRE, we leave this exploration for future work.
@@ 6.2.2. Overall FlashAttention Accuracy
To evaluate the accuracy of the FlashAttention attention on FSA, we compare it against torch.nn.functional.scaled_dot_product_attention in PyTorch (Ansel et al., 2024). We use the same head dimension and sequence lengths as in the performance evaluation. The input matrices $Q,K,V$ are randomly generated with a normal distribution in the range of $[0,1]$.
Table 2. Mean absolute error (MAE), mean squared error (MSE), and mean relative error (MRE) of the FlashAttention results on FSA.
SeqLen MAE MSE MRE
2048 1.269e-03 3.219e-04 2.532e-03
4096 3.022e-04 6.302e-05 6.043e-04
6144 8.742e-05 1.142e-08 1.747e-04
8192 1.234e-03 2.914e-04 2.470e-03
10240 9.175e-05 1.246e-08 1.832e-04
12288 9.071e-05 1.222e-08 1.811e-04
14336 2.563e-03 6.500e-04 5.125e-03
16384 1.021e-03 2.324e-04 2.041e-03
The overall FlashAttention accuracy results are shown in Table 2. The MAE lies in the range of $10^{-4}$ to $10^{-3}$, while the MRE is around $10^{-3}$ to $10^{-2}$. These results indicate that the piecewise linear approximation of the exp2 function has a negligible impact on the final attention results.
@@ 6.3. FSA Area Overhead
  • A Split unit within each processing element (PE), which separates input values into integer and fractional parts for the exp2 operation.
To evaluate the area overhead, we synthesize the systolic array portion of FSA, excluding SRAMs and DMA, at 1 GHz using a 16 nm commercial process. The resulting area breakdown is shown in Table 3.
@@
Standard PEs (excluding Split units) 84.10 16629356
Other logic 5.63 1111842
Total 89.73 17741199
FSA additional Upward data path 4.83 956556
Split units 4.78 945322
CMP units 0.65 128012
Total 10.26 2029891
As shown in Table 3, the standard systolic array occupies 89.73% of the total area, while FSA’s additional components contribute 10.26%. The dominant sources of overhead are the upward data path and the Split units, which account for 4.83% and 4.78% of the total area, respectively. This is due to their replication across all PEs in the $128\times 128$ array. In contrast, the CMP units contribute only 0.65%, as they are instantiated solely in the top row of the array.
@@ 7. Discussion
7.1. Computing Non-Linear Functions on Systolic Arrays
We observe that the piecewise linear approximation of the exp2 function on systolic arrays has strong potential to generalize to other non-linear functions.
In general, if a non-linear function can be segmented into multiple intervals, each well-approximated by a linear function, then piecewise linear approximation can be applied. This enables the computation of non-linear functions using the high-throughput FLOPs/s capacity of systolic arrays—substantially exceeding that of traditional vector units.
@@ 7.2. Necessity of a Vector Unit
While FSA is capable of executing FlashAttention entirely on the systolic array without relying on a vector unit, a vector or special function unit (SFU) is still necessary in a full accelerator system to support activation functions outside of FlashAttention.
However, because FSA offloads all attention-related computation from the vector unit, the overall vector FLOPs/s requirement may be significantly reduced.
@@ 7.3. FSA Variant: Area-Optimized Design
This results in the inner loop latency increasing from $105N+$ 10 to $106N+$ 10 cycles. Despite this, this area-optimized design would still deliver higher FLOPs/s utilization than commercial accelerators.
@@ 8. Conclusion
Beyond demonstrating the feasibility of executing complete FlashAttention computation on a single systolic array, our work establishes new possibilities for systolic array architectures to efficiently handle non-linear functions and complex operations that were previously deemed unsuitable for such architectures.

### v3

@@ 3.5. Overlapping Computations with SystolicAttention
During the first matrix multiplication, the head dimension $d$ of both $Q$ and $K$ is mapped to N_ROWSN\_ROWS, while the sequence length $B_{r}$ of $Q$ is mapped to N_COLSN\_COLS. After computing $S$, we re-stream it from the top of the array, mapping the $B_{c}$ dimension of $S$ to N_ROWSN\_ROWS. A sequence of element-wise operations is then performed in-place cycle by cycle to transform $S$ into $N$ and then into $P$.
For the second matrix multiplication, $O=PV$, the $B_{c}$ dimension of $V$ is mapped to N_ROWSN\_ROWS and contracted with the stationary matrix $P$ inside the systolic array.
@@
Figure 7 illustrates the execution of the entire inner loop. The fifth operation, a=old_mnew_ma=old\_m-new\_m, passes down the difference between the previous and current maximum values from comparison units to the accumulator. This operation runs in parallel with the constant multiplication $\frac{\log_{2}e}{\sqrt{d}}\cdot N$, as the downward data path is idle during the in-place constant multiplication.
Overall, each iteration takes only $2\times N\_COLS+3\times N\_ROWS+10$ cycles, meaning that on an $N\times N$ systolic array, it takes just $5N+10$ cycles to process a $N\times N$ FlashAttention tile. By contrast, on a naive $N\times N$ systolic array, the two independent matrix multiplications alone may require up to 8N28N-2 cycles due to preloading and synchronization overheads discussed in subsection 2.2.
The re-scaling operation in algorithm 1 of algorithm 1 is only performed once per outer loop. It is executed using the accumulator after the second matrix multiplication completes. In our implementation, this re-scaling step takes $2N+20$ cycles, which is negligible compared to the inner loop execution time.
@@ 4.1. Microarchitecture Overview
The instruction decoder classifies instructions into three types: load, store, and compute. Instructions of different classes are executed asynchronously, while instructions of the same class are issued in order but their execution may be interleaved.
@@ 5.2. Python API for FSA Instructions
Each FSA instruction is exposed through a corresponding Python API. These APIs are designed to be type-safe, ensuring correct usage of tensor types. The complete set of supported APIs is shown in 1.
@@ 5.3. JIT Kernel Compiler
The FlashAttention kernel implemented in FSA using the JIT compiler is shown in 2.
@@
In the example above, the host loads input tensors into device memory using DRAMSim2 (Rosenfeld et al., 2011), dispatches the compiled instructions to the FSA device, and retrieves the output tensors upon completion. Because FSA does not currently support hardware matrix transposition, the input matrix $V$ is transposed in advance, and the output matrix $O$ is converted to NumPy (Harris et al., 2020) arrays for post-processing. On commercial hardware, such transpositions can be performed by the DMA engine during memory transfers (AWS Neuron Team, 2025), and therefore should not incur significant overhead.
@@ 6.1. FlashAttention Performance
To clarify, we use the term FLOPs to denote the number of floating-point operations, and FLOPs/s to indicate the number of floating-point operations performed per second. Although all accelerators share the same systolic array size, their absolute performance still varies due to differences in clock frequency, which in turn depend on the specific node technology employed. Therefore, we simulate FSA at 1.5 GHz, a target that is achievable in 16 nm technology, as discussed in subsection 6.3, and compare matrix multiplication FLOPs/s utilization rather than wall-clock execution time. This metric reflects the end-to-end absolute performance of each accelerator under the same clock frequency.
@@
Systolic array size 128128 128128 128128
@@
Frequency 1.5GHz1 2.8GHz 1.5GHz
Memory Bandwidth 819GB/s 820GB/s 820GB/s
@@
The number of floating-point operations in FlashAttention (Dao, 2023) is given by:
$\text{Total FLOPs}=4\times\text{SeqLen}^{2}\times d$
where SeqLen is the sequence length and $d$ is the head dimension.
@@
$\text{Achieved FLOPs/s{}}=\frac{\text{Total FLOPs}}{\text{Execution Time}}$
@@
$\text{FLOPs/s{} Utilization}=\frac{\text{Achieved FLOPs/s{}}}{\text{Peak FLOPs/s{}}}$
@@
  • jax.experimental.pallas.ops.tpu.flash_attention for the TPUv5e (JAX Developers, [n. d.]).
@@
The experimental results are shown in Figure 11. Benefiting from intensively overlapped computations, FSA achieves average FLOPs/s utilization that is 1.77 times and 4.83 times higher compared to TPUv5e and AWS NeuronCore-v2, respectively. Moreover, the results confirm that our RTL implementation of FSA closely aligns with the theoretical performance outlined in subsection 3.5, validating the correctness of our RTL design.
@@ 6.2.1. Error Analysis of Piecewise Linear Approximation
Figure 12 shows the mean absolute error (MAE) and mean relative error (MRE) of the piecewise linear approximation of the exp2 function, plotted against the number of segments used in the approximation. The MAE decreases significantly as the number of segments increases, while the MRE remains relatively stable. This occurs because the breakpoints in the approximation are uniformly spaced, whereas the fp16 representation is logarithmic in nature, making relative error more sensitive for values close to zero. We choose to use 8 segments in our FSA implementation, which achieves a MAE of 0.00014 and an MRE of 0.02728. Although non-uniformly spaced breakpoints could further reduce the MRE, we leave this exploration for future work.
@@ 6.2.2. Overall FlashAttention Accuracy
To evaluate the accuracy of the FlashAttention attention on FSA, we compare it against torch.nn.functional.scaled_dot_product_attention in PyTorch (Ansel et al., 2024). We use the same head dimension and sequence lengths as in the performance evaluation. The input matrices are randomly generated using the same distribution as in the FlashAttention-3 (Shah et al., 2024) accuracy evaluation:
$Q,K,V\sim\mathcal{N}(0,1)+\mathcal{N}(0,100)\cdot\mathrm{Bernoulli}(0.001).$
Table 2. Mean absolute error (MAE), root mean squared error (RMSE), and mean relative error (MRE) of the FlashAttention results on FSA.
SeqLen MAE RMSE MRE
2048 7.983e-03 1.315e-02 1.558e-02
4096 1.379e-02 2.290e-02 2.596e-02
6144 1.849e-02 3.085e-02 3.545e-02
8192 2.253e-02 3.772e-02 4.413e-02
10240 2.595e-02 4.373e-02 5.259e-02
12288 2.890e-02 4.873e-02 5.920e-02
14336 3.165e-02 5.351e-02 6.529e-02
16384 3.403e-02 5.784e-02 7.181e-02
The overall accuracy results are summarized in Table 2. The MAE lies in the range of $7\times 10^{-3}$ to $4\times 10^{-2}$, while the MRE ranges from $1\times 10^{-2}$ to $8\times 10^{-2}$. These results indicate that the piecewise linear approximation of the exp2 function has a negligible impact on the final attention output.
@@ 6.3. FSA Area Overhead
  • A Split unit within each PE, which separates input values into integer and fractional parts for the exp2 operation.
To evaluate the area overhead, we synthesize the systolic array portion of FSA, excluding SRAMs and DMA, at 1.5 GHz using a 16 nm commercial process. The resulting area breakdown is shown in Table 3.
@@
Standard PEs 86.81 24445044
Other logic 1.11 313457
Total 87.92 24758501
FSA additional Upward data path 6.24 1756641
Split units 5.30 1493150
CMP units 0.53 149524
Total 12.07 2029891
As shown in Table 3, the standard systolic array occupies 87.92% of the total area, while FSA’s additional components contribute 12.07%. The dominant sources of overhead are the upward data path and the Split units, which account for 6.24% and 5.30% of the total area, respectively. This is due to their replication across all PEs in the $128\times 128$ array. In contrast, the CMP units contribute only 0.53%, as they are instantiated solely in the top row of the array.
7. Related Work
Significant efforts have been made to fuse FlashAttention on hardware accelerators. COSA (Wang et al., 2023) and COSA Plus (Wang et al., 2025) fuse FlashAttention across two cooperative systolic arrays, but data from the first array must still pass through a special function unit (SFU) for softmax. Matching SFU throughput to large systolic arrays can increase hardware cost. FuseMax (Nayak et al., 2024) builds on the cascade Einsum abstraction introduced in TeAAL (Nayak et al., 2023) to describe the FlashAttention algorithm and map it onto spatial accelerator models. FuseMax uses a different dataflow from the one used in our paper: by employing an output-stationary dataflow, it overlaps four FlashAttention iterations on the systolic array to obtain very high utilization. The overlapping of iterations necessitates storing multiple input tiles in SRAM and maintaining intermediate results within the array, leading to higher storage overhead and control complexity. Additionally, FuseMax assumes all FlashAttention operations are performed in fp16 format. In contrast, the original FlashAttention algorithm uses fp16 for matrix multiplications but fp32 for accumulations. Reconciling this difference may require intermediate results to be stored in fp32, which can further increase register usage. ExpMul (Alexandridis et al., 2025) fuses the exp and the second matrix multiplication in FlashAttention, which is orthogonal to our work and could be integrated into FSA as well.
@@ 8.1. Necessity of a Vector Unit
While FSA is capable of executing FlashAttention entirely on the systolic array without relying on a vector unit, a vector unit or SFU is still necessary in a full accelerator system to support activation functions outside of FlashAttention. However, because FSA offloads all attention-related computation from the vector unit, the overall vector FLOPs/s requirement may be significantly reduced.
@@ 8.2. FSA Variant: Area-Optimized Design
This results in the inner loop latency increasing from $5N+10$ to $6N+10$ cycles. Despite this, this area-optimized design would still deliver higher FLOPs/s utilization than commercial accelerators.